12. Stay Corporation uses the weighted-average method in its process costing system. The company's ending work in process inventory consists of 8,000 units, 60% complete with respect to materials and 80% complete with respect to conversion costs. If the total cost in this inventory is $200,000 and if the cost for materials is $16 per equivalent unit for the period, the conversion cost per equivalent unit of the production for the period must be:
Answer #1

Equivalent units of material = 8000*60% = 4800

Cost of material = 4800*16 = $76800

Equivalent units of conversion = 8000*80% = 6400

Cost of conversion = $200000-76800 = $123200

Conversion cost per equivalent unit = 123200/6400 = $19.25 per unit